OpenStudy (josh0404):

Can you please help me, here is the problem? The revenue R denied from selling x units of a commodity is computed as the product of x and the price P per unit. a. If the price demand equation is p=1600-10x, where P is the price (in pesos) per units when x units are sold, find the revenue when 60 units are sold. b. what is the maximum revenue in (a)? c. if C is the cost, in pesos, of producing x units of the commodity then, assuming that every unit produced is sold, the profit P is given by P=R-C, revenue minus cost, if the cost of producing x units of commodity in (a) is C=975+60x, what is

OpenStudy (josh0404):

help me please :( ?

OpenStudy (josh0404):

please help me

OpenStudy (josh0404):

thanks :)

MARC (marc_d):

a. substitute x=60 into the eqn p=1600-10x

MARC (marc_d):

\[p=1600-10(60)\]

OpenStudy (josh0404):

so it will become p= 1600-600

MARC (marc_d):

yep,then subtract 1600 with 600.

OpenStudy (josh0404):

the value of P=1000

MARC (marc_d):

yeap,that's right!

OpenStudy (josh0404):

how about the problem b and c Sir?

OpenStudy (josh0404):

what is the maximum revenue in a?

OpenStudy (josh0404):

Can you please help me, here is the problem? The revenue R denied from selling x units of a commodity is computed as the product of x and the price P per unit. a. If the price demand equation is p=1600-10x, where P is the price (in pesos) per units when x units are sold, find the revenue when 60 units are sold. b. what is the maximum revenue in (a)? c. if C is the cost, in pesos, of producing x units of the commodity then, assuming that every unit produced is sold, the profit P is given by P=R-C, revenue minus cost, if the cost of producing x units of commodity in (a) is C=975+60x, what is the profit when 60 units are produced and sold?

OpenStudy (josh0404):

Marc help me?

MARC (marc_d):

wait a minute,i think for question a,the question ask to find the revenue not the price.

OpenStudy (josh0404):

ok thank your time

MARC (marc_d):

np... equation for revenue \[R=p(x)\]

MARC (marc_d):

eqn for price \[p=1600-10x\]

OpenStudy (josh0404):

so what is the value of our Revenue?

MARC (marc_d):

substitute p=1600-10x into R=px

OpenStudy (josh0404):

what is the value of our x?

MARC (marc_d):

\[R=(1600-10x)(x)\]

MARC (marc_d):

in question a,it mention that 60 units are sold.

MARC (marc_d):

so,x is 60

MARC (marc_d):

sub. x=60 into the equation to find the revenue\[R=1600x-10x^2\]

OpenStudy (josh0404):

so our minimum Revenue is 60,000

MARC (marc_d):

yep,our revenue is 60000.

OpenStudy (josh0404):

how about the maximum Revenue how to solve it?

MARC (marc_d):

i'm x sure about question b but i know how to find for question c.

MARC (marc_d):

question c, sub. x=60 into this eqn C=975+60x

MARC (marc_d):

\[C=975+60(60)\]

MARC (marc_d):

\[C=4575\]

OpenStudy (josh0404):

show the solution marc, I have no idea how to solve it?

MARC (marc_d):

yea,that's what i'm doing right now... :)

MARC (marc_d):

okay for question C,they ask to find the profit when 60 units are produced and sold

MARC (marc_d):

given that the equation to find cost is\[C=975+60x\]

OpenStudy (josh0404):

yes Marc, then the formula for getting the profit is P=R-C?

MARC (marc_d):

yep,that's right

OpenStudy (josh0404):

so the profit is P=55,425 because our Revenue is 60,000-Cost 4575 then the total is 55,425

MARC (marc_d):

yes,that's correct! ^

OpenStudy (josh0404):

now lets proceed to problem b Marc?

OpenStudy (josh0404):

here is the example Marc i think this one help you to solve problem b

OpenStudy (acannell):

Can you hep me with my question?

MARC (marc_d):

it seems that question b needs to use differentiation method

OpenStudy (josh0404):

(1600-x)(975+60x) what do you think Marck?

OpenStudy (josh0404):

do you think my representation is right, to get the Maximum revenue?

MARC (marc_d):

i think it should be\[R=(1600-10x)(x)\]

MARC (marc_d):

it is based on question a) so,u need to use eqn for revenue and eqn for profit.

MARC (marc_d):

price*

OpenStudy (josh0404):

i think so

OpenStudy (josh0404):

show your proposal solution lets find out?

MARC (marc_d):

okay,here we go...\[R=1600x-10x^2\]

MARC (marc_d):

differentiate R

MARC (marc_d):

\[\frac{dR}{dx}=1600x^{1-1}-2(10)^{2-1}\]

OpenStudy (josh0404):

i think (Price)(cost) is the formula to get the maximum value, what do you think so?

MARC (marc_d):

we do not need to involve cost since the eqn for revenue is R=px

OpenStudy (josh0404):

a while ago we get the minimum value because the x is given at this case the x is unknown is in it?

MARC (marc_d):

for question a),it didn't say to get minimum value but revenue only. for question b),they ask to find the maximum revenue means that dR/dx=0

MARC (marc_d):

\[\frac{dR}{dx}=1600-20x\]

MARC (marc_d):

when maximum revenue,dR/dx=0

OpenStudy (josh0404):

where did you get your 20x?

OpenStudy (josh0404):

how is it, where did you get that solution?

MARC (marc_d):

i differentiate it\[\frac{dR}{dx}=1600x^{1-1}-2(10)x^{2-1}\]

OpenStudy (josh0404):

so what is the maximum value?

MARC (marc_d):

first we need to find the value of x(units for selling)\[\frac{dR}{dx}=1600-20x\]

MARC (marc_d):

when maximum revenue,dR/dx=0

MARC (marc_d):

\[0=1600-20x\]

MARC (marc_d):

find the value of x

MARC (marc_d):

20x=1600

MARC (marc_d):

x=80

OpenStudy (josh0404):

the maximum value is 80

MARC (marc_d):

not yet,that is just for the value of x( units sold)

MARC (marc_d):

now,do the samething as question a. Find the revenue using the value of x which is 80

OpenStudy (josh0404):

show the final answer, because i need to take a rest it already midnight in our country

MARC (marc_d):

yea,me too... xD Sorry,for taking ur time. R=px R=(price eqn)(x) R=(1600-10x)(x) R=(1600x-10x^2) R=1600(80)-10(80)^2

MARC (marc_d):

Maximum revenue,R=? solve it!

OpenStudy (josh0404):

thanks Marc :)

OpenStudy (josh0404):

i need to sleep now maybe we can continue our discussion tomorrow morning :D

OpenStudy (josh0404):

thanks for your Greatness i learn alot :)

OpenStudy (josh0404):

here bro

OpenStudy (josh0404):

dRdx=1600x1−1−2(10)x2−1 how did you get this formula bro?